OpenStudy (anonymous):

find the derivative of the trigonometric function…. f(theta)=(theta+1)cos theta

OpenStudy (anonymous):

f(theta)=(theta+1)cos theta

OpenStudy (anonymous):

is what I meant...

OpenStudy (anonymous):

use product formula

OpenStudy (anonymous):

that didn't help me understand how todo it…. I want to know step by step how to do it..

OpenStudy (anonymous):

cosQ(d/dx(Q-1))+(Q-1)(d/dx(cosQ)

OpenStudy (cwrw238):

product rule (f*g)' = f*g' + g*f ' f(x) = (x+1)cos x f\(x) = (x+1)*-sin x + 1* cos x = cos x - (1+x) sin x

OpenStudy (cwrw238):

for x read theta

OpenStudy (anonymous):

for x read theta?

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

yes i used x as the angle to replace theta thats all - as i'm too lazy to write theta... so answer is cos(theta) - (1 + theta ) sin (theta)

OpenStudy (anonymous):

ooohhhh ok ok !

OpenStudy (anonymous):

thankyou for explaining it to me. I really do appreciate your help. @cwrw238

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

would you use the product rule for y=x+cotx ?

OpenStudy (cwrw238):

no - this is a sum not a product just differentiate term by term the product rule is used where you have 2 functions multiplied together eg x^2 sin x , x(1 + tan x) , x^3 ln x

OpenStudy (anonymous):

ok I understand